OpenStudy (anonymous):

How do you find the root of an odd number? For example; 85squared.

OpenStudy (anonymous):

\[\sqrt{85^2}=85\] :D

OpenStudy (anonymous):

\[\sqrt{85}\] \[\sqrt{5 * 17}\] You can't for square root of 85. However, if it's some other number like for example, \[\sqrt{50}\] Then split it up like this using LCM, \[\sqrt{5*5*2}\] The final answer will be, \[5\sqrt{2}\]